I need help with this problem please:
Lane invested $14,000, part at 15% and part at 2%. If the total interest at the end of the year is $1,060, how much did she invest at 15%?
Thanks for your help!

Answer by checkley75(3666)

.15X+.02(14000-X)=1060
.15X+280-.02X=1060
.13X=1060-280
.13X=780
X=780/.13
X=6,000 INVESTED @ 15%
PROOF
.15*6000+.02*8000=1060
900+160=1060
1060=1060